# A Global Atlas of Digital Dermatology to Map Innovation and Disparities

### v1

@@ Appendix A Supplementary Figures
Supplementary Fig. 7: Pairwise Fréchet distance (FD) heatmap quantifying inter-dataset domain shifts. Matrix of FD between all 30 aggregated datasets. Dark blue indicates high similarity, statistically confirming subset relationships (e.g., Fitzpatrick17kSkinCAP) and institutional clusters (ISIC archives). Yellow/green bands identify high-uniqueness outliers that contribute significant distributional diversity to the global atlas.
@@ C.1.1 Sparse metadata against imputed
This structural consistency implies that the “missingness” of metadata in public archives is not driving the observed disparities. Rather, the bias is intrinsic to the image acquisition process itself. The unlabelled data essentially reflect the same exclusionary patterns as the labelled data: A heavy skew toward Western, middle-aged, and lightly pigmented populations. Consequently, the stark divergence from the global population statistics (right column) represents a genuine, systemic failure of the current dermatological data landscape to reflect the diversity of the human population.
@@ C.1.2 Missing data from a metadata perspective
While the univariate analysis presented in the main text exposes severe imbalances, such as the restriction of FST V–VI to only 5.8% of the archive and pediatric patients to 3.0%, these aggregate statistics potentially mask even deeper intersectional voids. To investigate the availability of data for specific patient subgroups, we analyzed the joint distributions of imputed metadata attributes (Supplementary Fig. 3). This intersectional audit reveals that the global archive is not merely skewed, but highly concentrated around a specific “standard” patient profile. As illustrated in the FST vs. age and gender vs. age heatmaps, the vast majority of the 1.1 million samples are clustered in a narrow demographic window: Patients aged 30–70 with FST II–III. This region of hypersaturation suggests that model performance reported in current literature is heavily weighted toward this specific demographic intersection. Conversely, the analysis highlights “dark corners” of the latent space where data support collapses entirely. Most critically, the intersection of age and skin tone reveals a compounded exclusion of pediatric populations with darker skin tones. While pediatric data is scarce overall, the bin representing pediatric patients (0–17 years) with FST V–VI is effectively empty. This implies that current foundation models possess virtually no ground-truth visual signals for common childhood dermatoses on darker skin, rendering them potentially unsafe for this vulnerable subgroup. Furthermore, the origin (continent) vs. age analysis underscores the geographical homogeneity of the data. Even when diversity in age or body region is present, it is almost exclusively derived from datasets originating from the Global North (specifically Europe and North America). The distinct lack of samples from the Global South (Africa and South America) across all body regions suggests that environmental and genetic variations specific to these populations are absent from the training dataset. Ultimately, these intersectional voids define the strict operational limits of the current Digital Atlas. Outside the saturated center of Western, middle-aged, fair-skinned patients, the reliability of diagnostic AI remains unsupported by the underlying data. The technical modality of data acquisition further compounds these structural exclusions (Supplementary Fig. 4). While the recent expansion of the atlas has been driven by the introduction of TBP, the intersectional analysis reveals that this modality is almost exclusively restricted to adult, fair-skinned populations (FST II–III) within Europe and North America. Specifically, TBP shows negligible coverage of pediatric patients (0.1%) and individuals with darker skin tones (FST V–VI), whereas it is heavily concentrated on the trunk and lower limbs.
@@ C.2.2 Dataset distances
To provide a comprehensive map of the relationships between the 30 aggregated datasets, we computed the pairwise FD between all collection pairs, visualized as a heatmap in Supplementary Fig. 7. This quantitative analysis reveals significant distributional homologies that challenge the assumption of dataset independence. We observe a distinct ISIC cluster where annual challenges (2016–2020) and institutional archives like MSKCC exhibit minimal pairwise distances, reflecting their shared lineage and patient populations. Furthermore, the analysis statistically confirms suspected subset relationships. The vanishingly small FD between Fitzpatrick17k and SkinCAP confirms that the latter functions as a subset of the former, while similar homologies link CAN2000 and CAN5600. Conversely, datasets such as PH2 and ISIC-challenge-2024 emerge as distributional outliers with high average distances to the rest of the field, identifying them as unique sources of variance and critical benchmarks for evaluating model robustness against genuine domain shifts.
@@ C.4 List of datasets
Our analysis integrates 30 publicly available dermatology image datasets, comprising 1,145,893 unique images after deduplication. Below, we provide a detailed description of each dataset, including imaging modality, geographic origin, and sample size. Dataset counts represent the final deduplicated samples retained in our corpus.
@@ C.4.2 Clinical photography datasets
  • MED-NODE [giotis2015a] (170 images): Computer-assisted melanoma diagnosis dataset from Netherlands, employing non-dermoscopic clinical photography.
@@ C.4.4 Educational and web-sourced datasets
  • DermaDia (9,848 images): Educational dermatology image collection of uncertain geographic provenance.

### v2

@@ Appendix A Supplementary Figures
Supplementary Fig. 7: Pairwise Fréchet distance (FD) heatmap quantifying inter-dataset domain shifts. Matrix of FD between all 29 aggregated datasets. Dark blue indicates high similarity, statistically confirming subset relationships (e.g., Fitzpatrick17kSkinCAP) and institutional clusters (ISIC archives). Yellow/green bands identify high-uniqueness outliers that contribute significant distributional diversity to the global atlas.
@@ C.1.1 Sparse metadata against imputed
This structural consistency implies that the “missingness” of metadata in public archives is not driving the observed disparities. Rather, the bias is intrinsic to the image acquisition process itself. The unlabelled data essentially reflect the same exclusionary patterns as the labelled data: A heavy skew toward Western, middle-aged, and lightly pigmented populations. Consequently, the stark divergence from global population statistics (right column) reflects a genuine, systemic failure of the current dermatological data landscape to capture the diversity of the human population.
@@ C.1.2 Missing data from a metadata perspective
While the univariate analysis presented in the main text exposes severe imbalances, such as the restriction of FST V–VI to only 5.8% of the archive and pediatric patients to 3.0%, these aggregate statistics potentially mask even deeper intersectional voids. To investigate the availability of data for specific patient subgroups, we analyzed the joint distributions of imputed metadata attributes (Supplementary Fig. 3). This intersectional audit reveals that the global archive is not merely skewed, but highly concentrated around a specific “standard” patient profile. As illustrated in the FST vs. age and gender vs. age heatmaps, the vast majority of the 1.1 million samples are clustered in a narrow demographic window: Patients aged 30–70 with FST II–III. This region of hypersaturation suggests that model performance reported in current literature is heavily weighted toward this specific demographic intersection. Conversely, the analysis highlights “dark corners” of the latent space where data support collapses entirely. Most critically, the intersection of age and skin tone reveals a compounded exclusion of pediatric populations with darker skin tones. While pediatric data is scarce overall, the bin representing pediatric patients (0–17 years) with FST V–VI is effectively empty. This implies that current foundation models possess virtually no ground-truth visual signals for common childhood dermatoses on darker skin, rendering them potentially unsafe for this vulnerable subgroup. Furthermore, the origin (continent) vs. age analysis underscores the geographical homogeneity of the data. Even when diversity in age or body region is present, it is almost exclusively derived from datasets originating from the Global North (specifically Europe and North America). The distinct lack of samples from the Global South (Africa and South America) across all body regions suggests that environmental and genetic variations specific to these populations are absent from the training dataset. Ultimately, these intersectional voids define the strict operational limits of the current Digital Atlas. Outside the saturated center of Western, middle-aged, fair-skinned patients, the reliability of diagnostic AI remains unsupported by the underlying data. The technical modality of data acquisition further compounds these structural exclusions (Supplementary Fig. 4). While the recent expansion of the atlas has been driven by the introduction of TBP, the intersectional analysis reveals that this modality is almost exclusively restricted to adult, fair-skinned populations (FST II–III) within Europe and North America. Specifically, TBP shows negligible coverage of pediatric patients and individuals with darker skin tones (FST V–VI), whereas it is heavily concentrated on the trunk and lower limbs.
@@ C.2.2 Dataset distances
To provide a comprehensive map of the relationships among the 29 aggregated datasets, we computed the pairwise FD between all collection pairs and visualized the results as a heatmap in Supplementary Fig. 7. This quantitative analysis reveals significant distributional homologies that challenge the assumption of dataset independence. We observe a distinct ISIC cluster in which annual challenges (2016–2020) and institutional archives such as MSKCC exhibit minimal pairwise distances, reflecting their shared lineage and patient populations. Furthermore, the analysis statistically confirms suspected subset relationships. The vanishingly small FD between Fitzpatrick17k and SkinCAP confirms that the latter functions as a subset of the former, while similar homologies link CAN2000 and CAN5600. Conversely, datasets such as PH2 and MED-NODE emerge as distributional outliers, with high average distances to the rest of the field, thereby identifying them as unique sources of variance and critical benchmarks for evaluating model robustness to genuine domain shifts.
@@ C.4 List of datasets
Our analysis integrates 29 publicly available dermatology image datasets, comprising 1,135,082 unique images after deduplication. Below, we provide a detailed description of each dataset, including imaging modality, geographic origin, and sample size. Dataset counts represent the final deduplicated samples retained in our corpus.
@@ C.4.2 Clinical photography datasets
  • MED-NODE [giotis2015a] (170 images): Computer-assisted melanoma diagnosis dataset from the Netherlands, employing non-dermoscopic clinical photography.